# Pay-per-search models are abstention models

### v1

@@ 3.1 Evaluation Details for Inference Mode I: w/ Search Tools
We compare MASH’s help-seeking model against the following baselines that also conduct RL training, but with different setups: (1) R1 trained using RL but without access to any search tools during training or evaluation. This baseline provides an upper bound for answer accuracy using only parametric knowledge. (2) Search-R1 (Jin et al., 2025) trained with search tools and a binary correctness reward; showcasing an upper bound without any penalties for searching, (3) OTC (Wang et al., 2025a) RL-trained for efficient search tool use. We compare these baselines to three MASH variants that differ in reward penalties (refer to § 2.2.2). Note that MASH w/ OTC and OTC differ in the warm-start procedure applied to the former.
Evaluation Metrics  We want our help-seeking model to strike a balance between answering parametrically (without search calls) and seeking help (with search calls). We report three metrics that collectively capture this: (1) Accuracy (Acc), i.e. if the predicted answer matches the gold response. Due to the limitations of exact match, we use an LLM judge, namely DeepSeek-V3.1 (Liu et al., 2024), to determine this. (2) Tool calls (TC), i.e. the average number of searches across trajectories. (3) Tool Productivity (TP) (Wang et al., 2025a), which is defined as $\sum_{i=1}^{|\mathcal{D}|}\nicefrac{{\mathbb{I}\{y_{i}=\hat{y}_{i}\}}}{{(1+m_{i})}}]/|\mathcal{D}|$ for test set $\mathcal{D}$. This discounts the accuracy of each output trajectory by its number of searches $m_{i}$. For all models, we report these metric averages over $4$ samples. We use TP on the validation set to select our model checkpoints for all methods, except Search-R1 for which we use accuracy; TP will result in a much inferior checkpoint selection for this case.
3.2 Evaluation Details for Inference Mode II: Abstention
In this evaluation mode, we follow the MASH process outlined in Figure 1 and § 2.1 to extract abstentions from a help-seeking model by removing access to search tools at inference.
Baselines  We compare against the following abstention baselines:
@@ 3.2 Evaluation Details for Inference Mode II: Abstention
Each of (1), (2) and (4) requires a definition of answerability; i.e. when can we claim that a question is answerable. A standard technique is to estimate the accuracy over $10$ samples and use a threshold $\lambda$ to classify into answerable or not. However, there does not exist a consensus in prior works on how to decide this threshold (Yang et al., 2024; Cheng et al., 2024). In our paper, we follow Yang et al. (2024) and set $\lambda=0.1$. Exact data curation and training details are in Appendix B.
Evaluation Metrics For abstention evaluation, we report two kinds of metrics:
  1. 1.
    Answer Accuracy: We report overall accuracy, i.e. over the test set, and precision, i.e. over non-abstained questions. Note that over-conservativeness, i.e. aggressively abstaining, will hurt overall accuracy but help precision, while under-conservativeness will have the opposite effect.
  2. 2.
    Abstention Classification: This captures whether a model’s abstention behavior is aligned with its knowledge boundaries, agnostic of accuracy. To avoid defining answerability (different reward penalties assume a different threshold), we evaluate over two groups of questions unaffected by the choice of $\lambda$: questions that the base models always answer incorrectly or always correctly. Let $\text{\%Abs}(0)$ and $\text{\%Abs}(1)$ be the percentage of questions for which a model abstains for the above two groups, respectively. We report $\%\text{Abs}(0)$ and Delta ($\%\text{Abs}(0)-\%\text{Abs}(1)$). A model recognizing its knowledge boundaries should have a high abstention rate for always incorrect questions, i.e. $\text{\%Abs}(0)$, and a much lower rate for always correct questions, captured by a large margin $\%\text{Abs}(0)-\%\text{Abs}(1)$. We do not evaluate the 2Wiki dataset for abstention classification due to there being only $58$ test examples in the $\text{Abs}(1)$ bucket, preventing reliable conclusions.
@@ 4.1 Inference Mode I: w/ Search Tools
Method Natural Questions HotPotQA 2Wiki
[…]
R1 26.04 0.0 26.04 26.53 0.0 26.53 9.18 0.0 9.18
Search-R1 (Jin et al., 2025) 57.31 1.0 28.66 56.34 3.00 14.09 45.39 3.00 11.35
MASH w/ otc 59.81 1.0 29.95 55.42 1.14 32.91 45.99 1.6 18.87
MASH w/ otc-st 56.37 0.64 38.63 53.32 1.10 32.55 46.23 1.64 19.08
MASH w/ exp 54.35 0.65 36.61 53.78 1.07 32.09 44.32 1.53 18.10
Table 1: Accuracy, average number of tool calls (TC) and tool productivity (TP) statistics for baselines and MASH evaluated under inference w/ search tools. MASH w/ OTC-ST is our best model with a $4.22\%$ and $5.62\%$ mean improvement on Acc and TP resp. over baseline OTC across datasets.
MASH outperforms all search baselines on tool productivity by effectively balancing accuracy and searches. Our results in Table 1 show that MASH, particularly MASH w/ OTC-Strict, leads to a $5.62$ point improvement on tool productivity over baseline OTC on average across datasets. Surprisingly, MASH variants report accuracies on par with Search-R1 (trained without any tool use penalty) on multi-hop datasets HotPotQA and 2Wiki, but with a substantially lower number of searches (e.g. $1.10$ and $1.64$ vs $3$). Moreover, this performance is a massive improvement over baseline OTC ($8.56\%$ and $6.64\%$ improvements on HotPotQA and 2Wiki respectively) with only a slightly higher number of searches. Tool productivity, which accounts for both these metrics, improves by 3.83 points on average over baseline OTC. Together, these results suggest that MASH not only reduces the number of searches, but also better operationalizes them to maintain accuracy.
Severe search penalties are needed for parametric answers for single-hop NaturalQA. We observed that both baseline OTC and MASH with the lenient OTC penalty (MASH w/ OTC) do not learn to answer parametrically for NaturalQA, i.e. converge to TC=1 for all questions. On the other hand, MASH w/ OTC-Strict answers parametrically for $36\%$ of the questions with only a $2.5\%$ drop in accuracy, thereby improving tool productivity by $9$ points. Similarly, MASH w/ Exp-Dec answers parametrically $35\%$, with a $4.5\%$ drop in accuracy [^3] Note that MASH w/ Exp-Dec training did result in checkpoints with higher accuracies. However, we use tool productivity on the validation set as the metric to select the final checkpoint for all methods. compared to baseline OTC but a 7 point improvement in tool productivity.
MASH variants extract better and more diverse search behaviors for multi-hop datasets via RL. Comparing search statistics for MASH w/ OTC and baseline OTC in Table 2, we see that they report a comparable number of parametric answers ($23.5\%$ vs $19.5\%$) but show very different search behaviors for the remaining questions. Particularly, the baseline OTC model without warm-start collapses to only one search for the remaining $80.2\%$ of its trajectories, while the warm-started model (MASH w/ OTC) can perform a mixture of one and multi-hop searches. In fact, MASH variants report a much higher accuracy for one search questions (56.456.4% vs 40.040.0%) by offloading the more “difficult” questions, i.e. those the model cannot answer with only one search, to the two search bucket. Baseline OTC fails to do this and reports lower overall accuracy. We see similar trends for the other multi-hop dataset, 2Wiki, as well.
@@
Table 2: Fine-grained tool use distribution (TC=0/1/2+ search) for baseline OTC and MASH models. We also report answer accuracies for questions in each subset (subscript). TC=0 indicates that the model answers parametrically. MASH can successfully off-load questions to parametric answering (from TC=1 to TC=0) will minimal or no decrease in accuracy (HotPotQA & NaturalQA).
MASH successfully aligns search tool use with parametric knowledge. For NaturalQA, the fine-grained search statistics in Table 2 show that the the questions that MASH w/ OTC-Strict and w/ Exp answer parametrically have similar answer accuracy compared to those for which they invoke one search call ($57.4$ vs $55.8$ for w/ OTC-Strict). This clearly shows that MASH can distinguish between parametrically answerable and not answerable questions and preferentially invoke tool calling for the latter to maintain overall accuracy.
@@ 4.2 Inference Mode II: w/ Abstention
MASH shows strong abstention behavior off-the-shelf. Table 3 (left) reports the answer accuracy for the overall test dataset (Acc) and the non-abstained questions (Prec) for each method. [^4] Note that it is possible to game one of these metrics by being over- or under-conservative. Therefore, all our conclusions are based on analyzing the two metrics collectively. First, we observe that, apart from MASH w/ OTC on NaturalQA, all MASH variants substantially outperform the prompting and Alignment for Honestly based SFT approaches in terms of answer precision and report comparable overall accuracy. In a couple of instances, we find that the AFH (Absolute) baseline reports better accuracy (e.g. HotPotQA and NaturalQA) compared to MASH, but this is accompanied by a 10-20% drop in precision.
We find that MASH w/ OTC-Strict, our best performing model from Section 4.1, is comparable to DPO for NaturalQA and HotPotQA; it outperforms DPO based on Prec. ($59.9$ vs $53.1$ for HotPotQA) but reports lower overall accuracy ($17.3$ vs $19.9$). We attribute this to the fact that MASH w/ OTC-Strict is more conservative, i.e. more likely to abstain, than DPO. For 2Wiki, MASH w/ OTC-Strict outperforms DPO on both Acc and Prec metrics.
@@
Table 3: Abstention accuracy (left) and abstention classification (rights) results for specialized abstention approaches and MASH. For abstention accuracy, we report Acc over the entire test set and Prec, i.e. accuracy over non-abstained answers. For classification, we report Abs(0), i.e. % abstention for unanswerable questions (higher better), and the delta (higher better) between the % abstention for unanswerable and answerable questions. We highlight runs with arguably similar performance but different trade-offs between Acc and Prec or Abs(0) and Delta in gray.
@@
5 Analysis
5.1 Analysis I: Impact of Warm-start on MASH Performance
@@
Table 4: MASH w/o warm-start tested in inference w/ search mode.
@@
5.2 Analysis II: Do oracle helpers improve selective help-seeking LLMs?
All experiments in Section 4 rely on a retrieval model (E5; Wang et al. (2022)) as the helper $H(\cdot)$. However, search results output by these retrievers can be noisy, which in turn generates a noisy signal for training the selective help-seeking LLM via RL. This prompts us to investigate if improving the “helper”, as opposed to the reward or initialization, can improve the learned help-seeking behavior.
Setup: We set $H(\cdot)$ to be an oracle; it directly returns the gold answer if the LLM invokes a help tag in its trajectory (exact prompts used is included in Appendix C). We train all MASH variants (OTC, OTC-Strict, Exp) for all datasets. Warm-start training is done for each individually with $l_{max}=1$.
Results: Help-seeking with oracle helpers fails to yield abstention behaviors. We find that every single training run converged to always asking for help within the first 50 training steps , even for the stricter help penalties. Note that the optimal policy should display selective help-seeking, i.e. answer parametrically for known questions, in order to maximize the chosen reward. However, we do not observe this in practice, as always seeking-help is an easy strategy for the LLMs to discover. For OTC and Exponential Decay, it is given non-zero rewards for all inputs. For OTC-Strict, it is given a positive reward for each question without correct parametric answers, which will be common early in training. This shows that the noisiness of the retrieval model is crucial to extract selective help-seeking over training, in a manner aligned with its parametric knowledge.
Figure 2: Average number of help requests for all MASH variants when trained with the oracle helper. All variants converge to 1 request within 20 steps.
Note that this setting with the oracle helper is equivalent to explicitly training for abstention using RL, with decreasing magnitude of rewards assigned for correct answers, abstention and incorrect answers. All training runs collapsing to always seeking help indicates that abstention training setting would also fail. We require RL algorithms with better exploration to succeed in this setting.
5.3 Analysis III: Out-of-Distribution Performance

### v2

@@ 3.1 Details for Inference Mode I: w/ Search Tools
We compare MASH’s help-seeking model against the following baselines that also conduct RL training, but with different setups: (1) R1 trained using RL but without access to any search tools during training or evaluation. This baseline provides an upper bound for answer accuracy using only parametric knowledge. (2) Search-R1 (Jin et al., 2025) trained with search tools and a binary correctness reward; showcasing an upper bound without any penalties for searching. We set the maximum number of searches for Search-R1 to 3 due to compute and memory concerns. (3) OTC (Wang et al., 2025a) RL-trained for efficient search tool use. We compare these baselines to three MASH variants that differ in reward penalties (refer to § 2.2.2). Note that MASH w/ OTC and OTC differ in the warm-start procedure applied to the former.
Evaluation Metrics  We want our help-seeking model to strike a balance between answering parametrically (without search calls) and seeking help (with search calls). We report three metrics that collectively capture this: (1) Accuracy (Acc), i.e. if the predicted answer matches the gold response. Due to the limitations of exact match, we use an LLM judge, namely DeepSeek-V3.1 (Liu et al., 2024), to determine this. (2) Tool calls (TC), i.e. the average number of searches across trajectories. (3) Tool Productivity (TP) (Wang et al., 2025a), which is defined as $\sum_{i=1}^{|\mathcal{D}|}\nicefrac{{\mathbb{I}\{y_{i}=\hat{y}_{i}\}}}{{(1+m_{i})}}]/|\mathcal{D}|$ for test set $\mathcal{D}$. This discounts the accuracy of each output trajectory by its number of searches $m_{i}$. For all models, we report these metric averages over $4$ samples. We use TP on the validation set to select our model checkpoints for all methods, except Search-R1 for which we use accuracy; TP results in a much inferior checkpoint selection for this case.
Table 1: Accuracy, average number of tool calls (TC) and tool productivity (TP) for baselines and MASH evaluated under inference w/ search tools. MASH w/ OTC-ST is our best model with a $4.22\%$ and $5.62\%$ mean improvement on Acc and TP over baseline OTC.
Method Natural Questions HotPotQA 2Wiki
[…]
R1 26.04 0.0 26.04 26.53 0.0 26.53 9.18 0.0 9.18
Search-R1 (Jin et al., 2025) 57.31 1.0 28.66 56.34 3.00 14.09 45.39 2.98 11.35
MASH w/ otc 59.81 1.0 29.95 55.42 1.14 32.91 45.99 1.6 18.87
MASH w/ otc-st 56.37 0.64 38.63 53.32 1.10 32.55 46.23 1.64 19.08
MASH w/ exp 54.35 0.65 36.61 53.78 1.07 32.09 44.32 1.53 18.10
3.2 Details for Inference Mode II: Abstention
In this mode, we follow the MASH process outlined in Figure 1 and § 2.1 to extract abstentions from a help-seeking model by removing access to search tools at inference.
Baselines  We compare against the following baselines:
@@ 3.2 Details for Inference Mode II: Abstention
  5. 5.
    Ternary, as considered by Xu et al. (2024a) and Wei et al. (2025), which conducts RL training with GRPO using a ternary reward assigning +1+1 for correct answers, 0 for abstention, and 1-1 abstention. We use the exact match reward from our other RL baselines to assess correctness.
Each of (1), (2) and (4) requires a definition of answerability; i.e. when we can claim that a question is answerable. A standard technique is to estimate the accuracy over $10$ samples and use a threshold $\eta$ to classify into answerable or not. However, there does not exist a consensus in prior works on how to decide this threshold (Yang et al., 2024; Cheng et al., 2024). In our paper, we follow (Yang et al., 2024) and set $\eta=0.1$. Exact data curation and training details are in Appendix C.
Evaluation Metrics For abstention evaluation, we report two kinds of metrics: (1) Answer Accuracy: We report overall Accuracy, i.e. over the test set, and Precision, i.e. over non-abstained questions. Note that over-conservativeness, i.e. aggressively abstaining, will hurt overall accuracy but help precision, while under-conservativeness will cause the opposite. (2) Abstention Classification: This captures whether a model’s abstention behavior is aligned with its knowledge boundaries, agnostic of accuracy. To avoid defining answerability (different reward penalties assume a different threshold), we evaluate over two groups of questions unaffected by the choice of $\eta$: questions that the base models always answer incorrectly, with an average accuracy of 0, or always correctly, with an average accuracy of $1$. Let $\text{\%Abs}(0)$ and $\text{\%Abs}(1)$ be the percentage of questions for which a model abstains for the above two groups, respectively. We report $\%\text{Abs}(0)$ and Delta ($\%\text{Abs}(0)-\%\text{Abs}(1)$). A model recognizing its knowledge boundaries should have a high abstention rate for always incorrect questions, i.e. $\text{\%Abs}(0)$, and a much lower rate for always correct questions, captured by a large margin $\%\text{Abs}(0)-\%\text{Abs}(1)$. We do not evaluate on 2Wiki for abstention classification as there are only $58$ test examples in the $\text{Abs}(1)$ bucket, preventing reliable conclusions.
@@ 4.1 Inference Mode I: w/ Search Tools
Table 2: Fine-grained tool use distribution (TC=0/1/2+ search) for baseline OTC and MASH. We also report answer accuracies for questions in each subset (subscript). For instance, MASH w/ OTC-Strict on 2Wiki answers questions with 2+ searches $77.5\%$ of time, with an accuracy of $49.2\%$. MASH successfully off-loads questions to parametric answering (from TC=1 to TC=0) with minimal or no decrease in accuracy (HotPotQA & NaturalQA). Highlight intensity correlates with the proportion of questions answered with TC=0/1/2+.
@@
MASH outperforms all search baselines on tool productivity by effectively balancing accuracy and searches. Our results in Table 1 show that MASH, particularly MASH w/ OTC-Strict, leads to a $5.62$ point improvement on tool productivity over baseline OTC on average across datasets. Surprisingly, MASH variants report accuracies on par with Search-R1 (trained without any tool use penalty) on multi-hop datasets HotPotQA and 2Wiki, but with a substantially lower number of searches (e.g. $1.10/1.64$ vs $3/2.98$). Moreover, this performance is a massive improvement over baseline OTC ($8.56\%$ and $6.64\%$ improvements on HotPotQA and 2Wiki respectively) with only a slightly higher number of searches. Tool productivity, which accounts for both these metrics, improves by 3.83 points on average over baseline OTC. Together, these results suggest that MASH not only reduces the number of searches, but also better operationalizes them to maintain accuracy.
Severe search penalties are needed for parametric answers for single-hop NaturalQA. We observed that both baseline OTC and MASH with the lenient OTC penalty (MASH w/ OTC) do not learn to answer parametrically for NaturalQA, i.e. converge to TC=1 for all questions. On the other hand, MASH w/ OTC-Strict answers parametrically for $36\%$ of the questions with only a $2.5\%$ drop in accuracy, thereby improving tool productivity by $9$ points. Similarly, MASH w/ Exp-Dec answers parametrically $35\%$ of the time, with a $4.5\%$ drop in accuracy [^4] Note that MASH w/ Exp-Dec training did result in checkpoints with higher accuracies. However, we use tool productivity on the validation set as the metric to select the final checkpoint. compared to baseline OTC but a 7 point improvement in tool productivity.
MASH variants extract better and more diverse search behaviors for multi-hop datasets. Comparing search statistics for MASH w/ OTC and baseline OTC in Table 2, we see that they report a comparable number of parametric answers ($23.5\%$ vs $19.5\%$) but show very different search behaviors for the remaining questions. Particularly, the baseline OTC model without warm-start collapses to only one search for the remaining $80.2\%$ of its trajectories, while the warm-started model (MASH w/ OTC) can perform a mixture of one and multi-hop searches. In fact, MASH variants report a much higher accuracy for one search questions (56.456.4% vs 40.040.0%) by offloading the more “difficult” questions, i.e. those the model cannot answer with only one search, to the two search bucket. Baseline OTC fails to do this and reports lower overall accuracy. We see similar trends for the other multi-hop dataset, 2Wiki, as well.
MASH successfully aligns search tool use with parametric knowledge. For NaturalQA, the fine-grained search statistics in Table 2 show that the questions MASH w/ OTC-Strict and w/ Exp answer parametrically have similar answer accuracy compared to those for which they invoke one search call ($57.4$ vs $55.8$ for w/ OTC-Strict). This clearly shows that MASH can distinguish between parametrically answerable and not answerable questions and preferentially invoke tool calling for the latter to maintain overall accuracy.
@@ 4.2 Inference Mode II: w/ Abstention
MASH shows strong abstention behavior off-the-shelf. Table 3 (left) reports the answer accuracy for the overall test dataset (Acc) and the non-abstained questions (Prec) for each method. [^5] Note that it is possible to game one of these metrics by being over- or under-conservative. Therefore, all our conclusions are based on analyzing the two metrics collectively. First, we observe that, apart from MASH w/ OTC on NaturalQA, all MASH variants substantially outperform the prompting and Alignment for Honesty based SFT approaches in terms of answer precision and report comparable overall accuracy. In a couple of instances, we find that the AFH (Absolute) baseline reports better accuracy (e.g. HotPotQA and NaturalQA) compared to MASH, but this is accompanied by a 10-20% drop in precision.
We find that MASH w/ OTC-Strict, our best performing model from Section 4.1, is comparable to DPO for NaturalQA and HotPotQA; it outperforms DPO based on Prec. ($59.9$ vs $53.1$ for HotPotQA) but reports lower overall accuracy ($17.3$ vs $19.9$). We attribute this to MASH w/ OTC-Strict being more conservative (i.e. more likely to abstain) than DPO. For 2Wiki, MASH w/ OTC-Strict outperforms DPO on both Acc and Prec. This competitive performance is despite MASH using the noisier exact match as the reward signal (due to our compute constraints), while DPO uses a stronger and more reliable LLM judge (DeepSeek-V3.1).
Table 3: Answer accuracy and abstention classification results for specialized abstention approaches and MASH. For answer accuracy, we report Acc over the entire test set and Prec, the accuracy over non-abstained answers. For classification, we report Abs(0), the abstention rate for unanswerable questions (higher better), and Delta, the margin between abstention rates for unanswerable and answerable questions (higher better). We highlight runs with similar performance but different trade-offs between Acc and Prec or Abs(0) and Delta in gray.
@@
Ternary 0.0 0.0 0.0 100.0 0.0 100.0 0.0
Finally, the Ternary baseline is outperformed by all methods and converges to always abstaining within 25 training steps on all datasets. This reproduces observations by Wei et al. (2025), which similarly found convergence to abstaining for a vast majority of questions with exact match as the correctness signal. Using a stronger LLM judge alternative is outside the scope of this paper due to resource limitations.
@@
4.3 Generalization to Other Model Scales and Families
We repeat experiments on Qwen2.5-7B base and Qwen3-4B base to determine whether MASH’s success generalizes to different parameter scales and model families, respectively. Due to the cost of replicating all RL experiments, we restrict our attention to HotPotQA and conduct RL training only for the OTC baseline and for MASH w/ OTC-Strict, the best performing variant for Qwen2.5-3B base.
MASH continues to outperform baselines at different model scales and families. Under the inference setting with search enabled, the OTC baseline converges to searching once for both models. MASH w/ OTC-Strict, on the other hand, consistently performs a mixture of parametric answers, and one- and multi-hop searches. Furthermore, it achieves average improvements of $2.98\%$ and $9.59$ on accuracy and tool productivity. Similar trends persist for abstention. MASH w/ OTC-Strict outperforms prompting and Alignment for Honesty. MASH yet again displays competitive performance with DPO, with different accuracy-precision trade-offs. Particularly, Qwen2.5-7B-Base achieves superior precision ($60.67\%$ vs $48.25\%$), while Qwen3-4B-Base achieves superior accuracy ($20.96\%$ vs $16.72\%$) over DPO. Result tables can be found in Appendix D.
4.4 Impact of Warm-start on Performance
Table 4: MASH w/o warm-start tested with search enabled.
@@ 4.4 Impact of Warm-start on Performance
4.5 Do oracle helpers improve selective help-seeking?
All experiments in Section 4 rely on a retrieval model (E5; (Wang et al., 2022)) as the helper $H(\cdot)$. However, search results output by these retrievers can be noisy, which in turn generates a noisy signal for training the selective help-seeking LLM via RL. This prompts us to investigate if improving the “helper”, as opposed to the reward or initialization, can improve the learned help-seeking behavior.
Setup: We set $H(\cdot)$ to be an oracle; it directly returns the gold answer if the LLM invokes a help tag in its trajectory (exact prompts used is included in Appendix E). We train all MASH variants (OTC, OTC-Strict, Exp) for all datasets. Warm-start training is done for each individually with $l_{max}=1$.
Results: Training with oracle helpers fails to yield selective help-seeking. We find that each training run converged to always asking for help within the first 50 training steps, even with stricter penalties. Note that the optimal policy should display selective help-seeking (answer parametrically for known questions) to maximize reward. However, we do not observe this in practice, as always seeking help is an easy strategy to discover. For OTC and Exponential Decay, it is given non-zero rewards for all inputs. For OTC-Strict, it is given a positive reward for each question without correct parametric answers, which is be common early in training. As such, the noisiness of the retrieval model is crucial to extract selective help-seeking in a manner aligned with its parametric knowledge. We require RL algorithms with better exploration to succeed in this oracle setting.
[…]
Natural Questions
otc 2.09 54.35 99.04 8.32
MASH w/ otc-st $18.25$ 51.24 79.95 51.63
DPO 24.4 - 77.38 68.23
TriviaQA
otc 4.07 71.43 96.95 7.11
MASH w/ otc-st 30.52 67.61 77.53 51.55
DPO 41.6 - 71.71 66.23
4.6 Out-of-Distribution Performance